OpenStudy (anonymous):

HELP!!! Urgently!!! Find the upper and lower sums of f(x) = -2x + 4 on the interval [0, 2] partitioned into four subintervals of equal length.

OpenStudy (amistre64):

define the height of the rectangles at multiples of 2/4

OpenStudy (anonymous):

I got that! But I don't know how to continue!

OpenStudy (amistre64):

multily each value by the width ... 1/2

OpenStudy (anonymous):

I'm confused!

OpenStudy (amistre64):

-2(0/2) + 4 -2(1/2) + 4 -2(2/2) + 4 -2(3/2) + 4 -2(4/2) + 4 4 3 2 1 0 multiply by 1/2 gives us the area of the rectangles with those given heights 4/2 3/2 2/2 1/2 0/2 add the first 4, then add the last 4

OpenStudy (anonymous):

But what would be the upper sum? and the lower sum?

OpenStudy (amistre64):

as i posted .... 4/2 3/2 2/2 1/2 0/2 add the first 4, then add the last 4

OpenStudy (anonymous):

I got 9.

OpenStudy (amistre64):

4/2 3/2 2/2 1/2 ---- 10/2 = 5 3/2 2/2 1/2 0/2 ----- 6/2 = 3

OpenStudy (anonymous):

Yeah! I just realize that! Thank you so much! I really appreciate it! Thank you!

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

Another question, for this problem, "Estimate the area of the given function, f(x) = x2 + 1 using 4 inscribed rectangles on the interval [2, 4]," should I have to use the same method you just show me?

OpenStudy (anonymous):

I figured it out! Thanks!